OpenStudy (anonymous):

A fixed quantity of gas at 22∘C exhibits a pressure of 760torr and occupies a volume of 5.02L . 1.)Calculate the volume the gas will occupy if the pressure is increased to 1.89atm while the temperature is held constant. 2.)Calculate the volume the gas will occupy if the temperature is increased to 170∘C while the pressure is held constant.

OpenStudy (dtan5457):

Just a bunch of gas laws to use here. For the first question, you can used the combined gas law. \[\frac{ P1V1 }{ T1 }=\frac{ P2V2 }{ T2 }\]

OpenStudy (dtan5457):

Actually no, if the temp is held constant, use a simpler formula (boyles law) v1p1=v2p2

OpenStudy (dtan5457):

convert your pressure units to the same unit (either torrs or atm)

OpenStudy (dtan5457):

For the 2nd question, use charles law v1/t1=v2/t2 make sure your temperature units are in kelvins

OpenStudy (dtan5457):

were you able to solve these?